OpenStudy (anonymous):

Determine whether the power function is even odd or neither f(x)=1/5x-11

OpenStudy (anonymous):

exponent is 1 is one an even or an odd?

OpenStudy (anonymous):

Exponent is -11

OpenStudy (anonymous):

so it's like \(y=\frac{1}{5}x^{-11}\)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

alright :) so let's see do you know the rule? Replace x with -x then you'll have to compare it with the function f(x) if the result is f(-x)=f(x) this function is an even f(-x)=-f(x) odd if they are not equal to f(x) nor -f(x) then its neither odd nor even

OpenStudy (anonymous):

\(f(x)=1/5x^{-11}\) sub x=-x \(f(-x)=1/5(-x)^{-11}\) \(f(-x)=-1/5x^{-11}\) \(f(-x)=-f(x)\)

OpenStudy (anonymous):

so it will be?

OpenStudy (anonymous):

Neither???

OpenStudy (anonymous):

No no no odd

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you !😊

OpenStudy (anonymous):

no problem :)